OpenStudy (anonymous):

percent notation of 9/16 my answer 9/1600 am i way off?

hero (hero):

To change any number to percent, multiply that number by 100, then add the percent sign.

hero (hero):

Basically the converted percentage has to be equal to the original number. If they are not equal, then you have done something wrong.

hero (hero):

Example: Convert 9 to a percentage form: First multiply 9 by 100: 9 x 100 = 900 Then add the percent sign: 900% Thus 9 = 900%

hero (hero):

Now why does 9 = 900%? Because % = 1/100

hero (hero):

So 9 = 900% 9 = 900/100 9 = 9

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

56.25

hero (hero):

What does that refer to?

hero (hero):

In general, if you are given something in one form, you should leave it in that same form after manipulation unless you are specifically told to write it in another form. What I mean by that is, you have a fraction. And you can write the number in percentage form without changing the number to the decimal. Basically, what I mean is, just leave the result in fraction form.

hero (hero):

First multiply 9/16 by 100 \[\frac{9}{16} \times 100 = \frac{225}{4} \] And then after adding the percent symbol: \[\frac{225}{4} \%\]

hero (hero):

So \[\frac{9}{16} = \frac{225}{4} \%\]